# Perfect state transfer between real pure states

### v1

@@ 10 Complete bipartite graphs
Let $\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix},\mathbf{y}\in\mathbb{R}^{m+n}$, where $\mathbf{x}_{1}\in\mathbb{R}^{m}$ and $|\sigma_{\mathbf{x}}(M)|\geq 3$. Let $\mathbf{z}$ and $\mathbf{v}^{\pm}=\begin{bmatrix}\sqrt{n}{\bf{1}}_{m}\\ \pm\sqrt{m}{\bf{1}}_{n}\end{bmatrix}$ be eigenvectors for $A(K_{m,n})$ associated with 0 and ±$\pm\sqrt{mn}$ respectively. Then $\mathbf{x}$ and $\mathbf{y}$ admit adjacency perfect state transfer in $K_{m,n}$ if and only if $\mathbf{x}\in\operatorname{span}\{\mathbf{v}^{+},\mathbf{v}^{-},\mathbf{z}\}$, $\mathbf{x}\notin\operatorname{span}\mathcal{U}$ for any two-subset $\mathcal{U}$ of $\{\mathbf{v}^{+},\mathbf{v}^{-},\mathbf{z}\}$ and $\mathbf{y}=\mathbf{x}-2\begin{bmatrix}\frac{1}{m}({\bf{1}}^{\top}\mathbf{x}_{1}){\bf{1}}\\ \frac{1}{n}({\bf{1}}^{\top}\mathbf{x}_{2}){\bf{1}}\end{bmatrix}$. Moreover, the minimum PST time is $\frac{\pi}{\sqrt{mn}}$.
@@
Taking $\mathbf{y}=U_{A}(\frac{\pi}{\sqrt{mn}})$ yields the desired conclusion. ∎
@@
Note that $\mathbf{e}_{u}-\mathbf{e}_{w}$ is fixed whenever $u$ and $w$ are non-adjacent. Now, suppose $u$ and $w$ are adjacent. Then equation (11) yields $U_{A}(t)(\mathbf{e}_{u}-\mathbf{e}_{w})=\begin{bmatrix}\mathbf{e}_{u}-\frac{2}{m}{\bf{1}}\\ -(\mathbf{e}_{w}-\frac{2}{m}{\bf{1}})\end{bmatrix}$, and so (1) and (2) are immediate. ∎
@@
If $u$ and $w$ are non-adjacent, then the same argument in Corollary 53 proves statement (i). Otherwise, equation (11) yields $U_{A}(t)(\mathbf{e}_{u}-\mathbf{e}_{w})=\begin{bmatrix}\mathbf{e}_{u}+\mathbf{e}_{w}-\frac{4}{m}{\bf{1}}\\ {\bf{0}}\end{bmatrix}$, and so statement (ii) is immediate. ∎
@@
Let $\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix},\mathbf{y}\in\mathbb{R}^{m+n}$, where $\mathbf{x}_{1}\in\mathbb{R}^{m}$ and where $|\sigma_{\mathbf{x}}(M)|\geq 3$. Let $\mathbf{u}$, $\mathbf{v}$ and $\mathbf{w}$ be eigenvectors for $L(K_{m,n})$ associated with $m$, $n$ and $m+n$ respectively. Then $\mathbf{x}$ and $\mathbf{y}$ admit Laplacian perfect state transfer in $K_{m,n}$ if and only if $\mathbf{x}\in\operatorname{span}\{{\bf{1}},\mathbf{u},\mathbf{v},\mathbf{w}\}$, $\mathbf{x}\notin\operatorname{span}\mathcal{U}$ for any two-subset $\mathcal{U}$ of $\{{\bf{1}},\mathbf{u},\mathbf{v},\mathbf{w}\}$ and either
@@
    $\nu_{2}(m)<\nu_{2}(n)$ and $\mathbf{y}=\begin{bmatrix}\mathbf{x}_{1}+\frac{2}{m+n}(({\bf{1}}^{\top}\mathbf{x}_{2})-\frac{n}{m}({\bf{1}}^{\top}\mathbf{x}_{1})){\bf{1}}\\ -\mathbf{x}_{2}+\frac{2}{m+n}(({\bf{1}}^{\top}\mathbf{x}_{1})+({\bf{1}}^{\top}\mathbf{x}_{2}){\bf{1}}){\bf{1}}\end{bmatrix}$.
@@
If $u$ and $w$ are non-adjacent, then $\mathbf{e}_{u}-\mathbf{e}_{w}$ is fixed. Otherwise, $\mathbf{e}_{u}-\mathbf{e}_{w}$ has eigenvalue support $\{m,n,m+n\}$. Applying Theorem 55(1) with $\mathbf{x}_{1}=\mathbf{e}_{u}$ and $\mathbf{x}_{2}=-\mathbf{e}_{v}$ yields the desired conclusion. ∎
@@ 11 Minimizing PST time
    Amongs all connected unweighted $n$-vertex graphs, $\mathbf{x}$ attains the least minimum period in $G$ relative to $L$ if and only if $G=G_{1}\vee G_{2}$ with $|V(G_{i})|=n_{i}$ for $i\in\{1,2\}$, $n=n_{1}+n_{2},$ and $\mathbf{x}\in\operatorname{span}\left\{{\bf{1}}_{n},\begin{bmatrix}n_{2}{\bf{1}}_{n_{1}}\\ -n_{1}{\bf{1}}_{n_{2}}\end{bmatrix}\right\}$.
@@
Let $\lambda_{1}$ and $\lambda_{2}$ be the largest and smallest eigenvalues of $M$. We first prove 1. By assumption, 00 is a simple eigenvalue of $L(G)$ with eigenvector ${\bf{1}}_{n}$. Moreover, every eigenvalue $\lambda$ of $L(G)$ satisfies $\lambda\leq n$ with equality if and only if $G$ is a join graph. Thus, for any two eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of $L(G)$, the Laplacian spread $\lambda_{1}-\lambda_{2}$ is maximum if and only if $\lambda_{1}=n$ and $\lambda_{2}=0$. Invoking Lemma 10, the least minimum period is attained if and only if $G=G_{1}\vee G_{2}$ for some graphs $G_{i}$ on $n_{i}$ vertices, $i\in\{1,2\}$ and $\sigma_{\mathbf{x}}(L)=\{0,n\}$, in which case $\rho=\frac{\pi}{n}$ and $\begin{bmatrix}n_{2}{\bf{1}}_{n_{1}}\\ -n_{1}{\bf{1}}_{n_{2}}\end{bmatrix}$ is the eigenvector associated with $n$. To prove 2, we use a result due to Breen, Riasanovsky, Tait and Urschel [BRTU] states that there is an $N>0$ such that if $n\geq N,$ the maximum adjacency spread $\lambda_{1}-\lambda_{2}$ over all connected $n$-vertex graphs is attained uniquely by the complete split graph $G=O_{a}\vee K_{n-a}.$ In this case, we have $\lambda_{1}=\lambda^{+}$, $\lambda_{2}=\lambda^{-}$ and $\lambda^{\pm}=\frac{1}{2}(n-a-1\pm\sqrt{(n-a-1)^{2}+4a(n-a)})$ so that $\lambda_{1}-\lambda_{2}=\sqrt{(n-a-1)^{2}+4a(n-a)}\approx\frac{2n}{\sqrt{3}}.$ The same argument used in the above case yields the desired conclusion. ∎
@@
    There exists an integer $N>0$ such that amongst all connected unweighted $n$-vertex graphs with $n\geq N$, $\mathbf{x}$ and $\mathbf{y}$ attain the least minimum PST time relative to $A$ if and only if the conditions in Theorem 58(2) hold and $\mathbf{y}$ has the form below, where where $D^{\pm}=\pm a(\lambda^{\pm})\sqrt{(n-a-1)^{2}+4a(n-a)}$.
@@ 12 Sensitivity with respect to readout time
$\frac{d^{k}f}{dt^{k}}\Big{|}_{\tau}=\begin{cases}(-1)^{\frac{k\mod 4}{2}}\sum_{j=0}^{k}(-1)^{j}{{k}\choose{j}}\mathbf{y}^{\top}M^{j}\mathbf{y}\mathbf{y}^{\top}M^{k-j}\mathbf{y}&{\rm{if}}\ k\ \rm{is\ even}\\ 0&{\rm{if}}\ k\ \rm{is\ odd}.\end{cases}$
Invoking Lemma 18, we have $\mathbf{x}^{\top}M^{k}\mathbf{x}=\mathbf{y}^{\top}M^{k}\mathbf{y}$ for all integers $k\geq 0$
There is practical interest in determining how large or small $\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}$ can be, since that will give intuition on what features govern the sensitivity of the fidelity with respect to the readout time. Note that
$\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}=-2(\mathbf{y}^{\top}M^{2}\mathbf{y}+(\mathbf{y}^{\top}M\mathbf{y})^{2}).$ (12)
@@
$\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}=2(\mathbf{y}^{\top}M^{2}\mathbf{y}+(\mathbf{y}^{\top}M\mathbf{y})^{2})=2\left(\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}+(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}\right),$
where $a_{j}=c_{j}^{2}$ for each $j=1,\ldots,\ell.$ Let $j_{0}$ denote an index such that $|\lambda_{j_{0}}|=\max\{|\lambda_{j}||j=1,\ldots,\ell\}.$ It is now straightforward to determine that $\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}\geq-4\lambda_{j_{0}}^{2}.$ This lower bound may be approached arbitrarily closely when $\mathbf{x}$ is of the form $\epsilon\mathbf{u}+\sqrt{1-\epsilon^{2}}\mathbf{v},$ where $\mathbf{v}$ is an eigenvector corresponding to $\lambda_{j_{0}}$, $\mathbf{u}$ is an eigenvector corresponding to $\lambda_{j}$ for some $j\neq j_{0},$ and $|\epsilon|$ is small.
@@
Next we seek an attainable lower bound on $\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}+(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}$ subject to the constraint that $\sum_{j=1}^{\ell}a_{j}=1$ and $a_{j}\geq 0$ for all $j=1,\ldots,\ell.$ Suppose that we have distinct indices $j_{1},j_{2}$ such that $a_{j_{1}},a_{j_{2}}>0.$ For each $j=1,\ldots,\ell$, consider the coefficients $b_{j}$ given by $b_{j}=a_{j}$ whenever $j\neq j_{1},j_{2},b_{j_{1}}=a_{j_{1}}+\epsilon$, and $b_{j_{2}}=a_{j_{2}}-\epsilon,$ where $\epsilon$ is sufficiently small. It is straightforward to show that
$\begin{split}\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}+(\sum_{j=1}^{\ell}b_{j}\lambda_{j})^{2}&=\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}+\left(\sum_{j=1}^{\ell}a_{j}\lambda_{j}\right)^{2}+\epsilon(\lambda_{j_{1}}-\lambda_{j_{2}})\left(\lambda_{j_{1}}+\lambda_{j_{2}}+2\sum_{j=1}^{\ell}a_{j}\lambda_{j}\right)+\epsilon^{2}(\lambda_{j_{1}}-\lambda_{j_{2}})^{2}\end{split}$
@@
Hence, if $\lambda_{j_{1}}+\lambda_{j_{2}}+2\sum_{j=1}^{\ell}a_{j}\lambda_{j}\neq 0$, then we can choose a value of $\epsilon$so that $\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}+(\sum_{j=1}^{\ell}b_{j}\lambda_{j})^{2}<\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}+(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}.$ Now observe that if there is a third index $j_{3}\neq j_{1},j_{2}$ such that $a_{j_{3}}>0,$ then necessarily one of $\lambda_{j_{1}}+\lambda_{j_{2}}+2\sum_{j=1}^{\ell}a_{j}\lambda_{j}$ and $\lambda_{j_{1}}+\lambda_{j_{3}}+2\sum_{j=1}^{\ell}a_{j}\lambda_{j}$ is nonzero, so that $\lambda_{j_{1}}+\lambda_{j_{2}}+2\sum_{j=1}^{\ell}a_{j}\lambda_{j}$ cannot attain the minimum value. Consequently, in order to find the minimum value, it suffices to focus on expressions of the form $a\lambda_{j_{1}}^{2}+(1-a)\lambda_{j_{2}}^{2}+(a\lambda_{j_{1}}+(1-a)\lambda_{j_{2}})^{2}$ where $a\in[0,1].$
Considering this last expression as a function of $a,$ we see that it is quadratic in $a$. It follows that one of two cases applies: i) either the vertex of the parabola is in $[0,1]$ in which case it yields the minimizing values, or ii) the vertex falls outside of $[0,1]$, in which case the minimum value is taken at either $a=0$ or $a=1.$ It is straightforward to determine that the vertex of the parabola corresponds to $a=-\frac{\lambda_{j_{1}}+3\lambda_{j_{2}}}{2(\lambda_{j_{1}}-\lambda_{j_{2}})}.$ Assume without loss of generality that $\lambda_{j_{1}}>\lambda_{j_{2}}.$ We find that the vertex of the parabola falls in $(0,1)$ if and only if $03\lambda_{j_{1}}+\lambda_{j_{2}}>$ 0 and $\lambda_{j_{1}}+3\lambda_{j_{2}}<0.$ When those conditions hold, we find that the minimum value is given by $-\frac{1}{4}(\lambda_{j_{1}}^{2}+6\lambda_{j_{1}}\lambda_{j_{2}}+\lambda_{j_{2}}^{2}).$ This covers case i). We note in passing that if $03\lambda_{j_{1}}+\lambda_{j_{2}}>$ 0 and $\lambda_{j_{1}}>\lambda_{j_{2}},$ then necessarily $\lambda_{j_{1}}>0.$ Since $\lambda_{j_{1}}+3\lambda_{j_{2}}<0$ it must also be the case that $\lambda_{j_{2}}<0.$ If case ii) holds, we find readily that the minimum value is equal to $2\min\{\lambda_{j_{1}}^{2},\lambda_{j_{2}}^{2}\}.$
We summarize the above discussion as follows.
@@
Suppose perfect state transfer occurs between unit vectors $\mathbf{x}$ and $\mathbf{y}$ at time $\tau$ relative to $M$. Define the sets $B_{1}=\{-4\lambda_{j}^{2}\ |\ \lambda_{j}\in\sigma_{\mathbf{x}}(M)\}$ and
[…]
Then
$\max\{B_{1}\cup B_{2}\}\ \geq\ \frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}\ \geq\ \min B_{1}.$
@@
Consider the Petersen graph with adjacency matrix $A,$ which has eigenvalues $3,1,-2$. Let $\bf{v}_{1},\bf{v}_{2},\bf{v}_{3}$ be unit eigenvectors corresponding to $3,1$ and $-2,$ respectively and form $\mathbf{x}\in\mathbb{R}^{10}$ as $\mathbf{x}=c_{1}\bf{v}_{1}+c_{2}\bf{v}_{2}+c_{3}\bf{v}_{3}$ where $c_{1},c_{2},c_{3}\in\mathbb{R}\backslash\{0\},c_{1}^{2}+c_{2}^{2}+c_{3}^{2}=1.$ Setting $\mathbf{y}=-c_{1}\bf{v}_{1}-c_{2}\bf{v}_{2}+c_{3}\bf{v}_{3}$, we find that there is PST from $\mathbf{x}\mathbf{x}^{\top}$ to $\mathbf{y}\mathbf{y}^{\top}$ and time $\pi.$ According to Theorem 61, $\frac{d^{2}f}{dt^{2}}\Big{|}_{\tau}$ is bounded below by $-36$ and above by $-\frac{7}{2}.$ The latter corresponds to the choices $\lambda_{j_{1}}=1,\lambda_{j_{2}}=-2.$
Acknowledgement. C. Godsil is supported by NSERC grant number RGPIN-9439. S. Kirkland is supported by NSERC grant number RGPIN–2019–05408. H. Monterde is supported by the University of Manitoba Faculty of Science and Faculty of Graduate Studies.

### v2

@@ 10 Complete bipartite graphs
Let $\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix},\mathbf{y}\in\mathbb{R}^{m+n}$, where $\mathbf{x}_{1}\in\mathbb{R}^{m}$ and $|\sigma_{\mathbf{x}}(A)|\geq 3$. Let $\mathbf{z}$ and $\mathbf{v}^{\pm}=\begin{bmatrix}\sqrt{n}{\bf{1}}_{m}\\ \pm\sqrt{m}{\bf{1}}_{n}\end{bmatrix}$ be eigenvectors for $A(K_{m,n})$ associated with 0 and ±$\pm\sqrt{mn}$ respectively. Then $\mathbf{x}$ and $\mathbf{y}$ admit adjacency perfect state transfer in $K_{m,n}$ if and only if $\mathbf{x}\in\operatorname{span}\{\mathbf{v}^{+},\mathbf{v}^{-},\mathbf{z}\}$, $\mathbf{x}\notin\operatorname{span}\mathcal{U}$ for any two-subset $\mathcal{U}$ of $\{\mathbf{v}^{+},\mathbf{v}^{-},\mathbf{z}\}$ and $\mathbf{y}=\mathbf{x}-2\begin{bmatrix}\frac{1}{m}({\bf{1}}^{\top}\mathbf{x}_{1}){\bf{1}}\\ \frac{1}{n}({\bf{1}}^{\top}\mathbf{x}_{2}){\bf{1}}\end{bmatrix}$. Moreover, the minimum PST time is $\frac{\pi}{\sqrt{mn}}$.
@@
Taking $\mathbf{y}=U_{A}(\frac{\pi}{\sqrt{mn}})\mathbf{x}$ yields the desired conclusion. ∎
@@
Note that $\mathbf{e}_{u}-\mathbf{e}_{w}$ is fixed whenever $u$ and $w$ are non-adjacent. Now, suppose $u$ and $w$ are adjacent. Then equation (12) yields $U_{A}(t)(\mathbf{e}_{u}-\mathbf{e}_{w})=\begin{bmatrix}\mathbf{e}_{u}-\frac{2}{m}{\bf{1}}\\ -(\mathbf{e}_{w}-\frac{2}{n}{\bf{1}})\end{bmatrix}$, and so (1) and (2) are immediate. ∎
@@
If $u$ and $w$ are adjacent, then the same argument in Corollary 53 proves statement (i). Otherwise, equation (12) yields $U_{A}(t)(\mathbf{e}_{u}+\mathbf{e}_{w})=\begin{bmatrix}\mathbf{e}_{u}+\mathbf{e}_{w}-\frac{4}{m}{\bf{1}}\\ {\bf{0}}\end{bmatrix}$, and so statement (ii) is immediate. ∎
@@
Let $\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix},\mathbf{y}\in\mathbb{R}^{m+n}$, where $\mathbf{x}_{1}\in\mathbb{R}^{m}$ and where $|\sigma_{\mathbf{x}}(L)|\geq 3$. Let $\mathbf{u}$, $\mathbf{v}$ and $\mathbf{w}$ be eigenvectors for $L(K_{m,n})$ associated with $m$, $n$ and $m+n$ respectively. Then $\mathbf{x}$ and $\mathbf{y}$ admit Laplacian perfect state transfer in $K_{m,n}$ if and only if $\mathbf{x}\in\operatorname{span}\{{\bf{1}},\mathbf{u},\mathbf{v},\mathbf{w}\}$, $\mathbf{x}\notin\operatorname{span}\mathcal{U}$ for any two-subset $\mathcal{U}$ of $\{{\bf{1}},\mathbf{u},\mathbf{v},\mathbf{w}\}$ and either
@@
    $\nu_{2}(m)<\nu_{2}(n)$ and $\mathbf{y}=\begin{bmatrix}\mathbf{x}_{1}+\frac{2}{m+n}(({\bf{1}}^{\top}\mathbf{x}_{2})-\frac{n}{m}({\bf{1}}^{\top}\mathbf{x}_{1})){\bf{1}}\\ -\mathbf{x}_{2}+\frac{2}{m+n}(({\bf{1}}^{\top}\mathbf{x}_{1})+({\bf{1}}^{\top}\mathbf{x}_{2})){\bf{1}}\end{bmatrix}$.
@@
If $u$ and $w$ are non-adjacent, then $\mathbf{e}_{u}-\mathbf{e}_{w}$ is fixed. Otherwise, $\mathbf{e}_{u}-\mathbf{e}_{w}$ has eigenvalue support $\{m,n,m+n\}$. Applying Theorem 55(1) with $\mathbf{x}_{1}=\mathbf{e}_{u}$ and $\mathbf{x}_{2}=-\mathbf{e}_{w}$ yields the desired conclusion. ∎
@@ 11 Minimizing PST time
    Amongst all connected unweighted $n$-vertex graphs, $\mathbf{x}$ attains the least minimum period in $G$ relative to $L$ if and only if $G=G_{1}\vee G_{2}$ with $|V(G_{i})|=n_{i}$ for $i\in\{1,2\}$, $n=n_{1}+n_{2},$ and $\mathbf{x}\in\operatorname{span}\left\{{\bf{1}}_{n},\begin{bmatrix}n_{2}{\bf{1}}_{n_{1}}\\ -n_{1}{\bf{1}}_{n_{2}}\end{bmatrix}\right\}$.
@@
Let $\lambda_{1}$ and $\lambda_{2}$ be the largest and smallest eigenvalues of $M$. We first prove 1. By assumption, 00 is a simple eigenvalue of $L(G)$ with eigenvector ${\bf{1}}_{n}$. Moreover, every eigenvalue $\lambda$ of $L(G)$ satisfies $\lambda\leq n$ with equality if and only if $G$ is a join graph. Thus, for any two eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of $L(G)$, the Laplacian spread $\lambda_{1}-\lambda_{2}$ is maximum if and only if $\lambda_{1}=n$ and $\lambda_{2}=0$. Invoking Lemma 10, the least minimum period is attained if and only if $G=G_{1}\vee G_{2}$ for some graphs $G_{i}$ on $n_{i}$ vertices, $i\in\{1,2\}$ and $\sigma_{\mathbf{x}}(L)=\{0,n\}$, in which case $\rho=\frac{2\pi}{n}$ and $\begin{bmatrix}n_{2}{\bf{1}}_{n_{1}}\\ -n_{1}{\bf{1}}_{n_{2}}\end{bmatrix}$ is the eigenvector associated with $n$. To prove 2, we use a result due to Breen, Riasanovsky, Tait and Urschel [BRTU] states that there is an $N>0$ such that if $n\geq N,$ the maximum adjacency spread $\lambda_{1}-\lambda_{2}$ over all connected $n$-vertex graphs is attained uniquely by the complete split graph $G=O_{a}\vee K_{n-a}.$ In this case, we have $\lambda_{1}=\lambda^{+}$, $\lambda_{2}=\lambda^{-}$ and $\lambda^{\pm}=\frac{1}{2}(n-a-1\pm\sqrt{(n-a-1)^{2}+4a(n-a)})$ so that $\lambda_{1}-\lambda_{2}=\sqrt{(n-a-1)^{2}+4a(n-a)}\approx\frac{2n}{\sqrt{3}}.$ The same argument used in the above case yields the desired conclusion. ∎
@@
    There exists an integer $N>0$ such that amongst all connected unweighted $n$-vertex graphs with $n\geq N$, $\mathbf{x}$ and $\mathbf{y}$ attain the least minimum PST time relative to $A$ if and only if the conditions in Theorem 58(2) hold and $\mathbf{y}$ has the form below, where $D^{\pm}=\pm a(\lambda^{\pm})\sqrt{(n-a-1)^{2}+4a(n-a)}$.
@@ 12 Sensitivity with respect to readout time
$\frac{d^{k}f}{dt^{k}}\bigg{|}_{\tau}=\begin{cases}(-1)^{\frac{k\mod 4}{2}}\displaystyle\sum_{j=0}^{k}(-1)^{j}{{k}\choose{j}}(\mathbf{y}^{\top}M^{j}\mathbf{y})(\mathbf{y}^{\top}M^{k-j}\mathbf{y})&{\rm{if}}\ k\ \rm{is\ even}\\ 0&{\rm{if}}\ k\ \rm{is\ odd}.\end{cases}$ (13)
Moreover, by Lemma 18, we have $\mathbf{x}^{\top}M^{k}\mathbf{x}=\mathbf{y}^{\top}M^{k}\mathbf{y}$ for all integers $k\geq 0$.
There is practical interest in determining how large or small $\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}$ can be, since that will give intuition on what features govern the sensitivity of the fidelity with respect to the readout time. Taking $k=2$ in (13) yields
$\begin{split}\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}=-2(\mathbf{y}^{\top}M^{2}\mathbf{y}-(\mathbf{y}^{\top}M\mathbf{y})^{2}).\end{split}$ (14)
@@
$\frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}=2((\mathbf{y}^{\top}M\mathbf{y})^{2}-\mathbf{y}^{\top}M^{2}\mathbf{y})=2((\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}),$
where $a_{j}=c_{j}^{2}$ for each $j=1,\ldots,\ell.$ Observe that from the Cauchy-Schwarz inequality,
$(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}=(\sum_{j=1}^{\ell}(c_{j}\lambda_{j})c_{j})^{2}\leq(\sum_{j=1}^{\ell}c_{j}^{2}\lambda_{j}^{2})(\sum_{j=1}^{\ell}c_{j}^{2})=\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}.$
We thus deduce that $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}\leq 0.$ This upper bound may be approached arbitrarily closely when $\mathbf{x}$ is of the form $\epsilon\mathbf{u}+\sqrt{1-\epsilon^{2}}\mathbf{v},$ where $\mathbf{u}$ and $\mathbf{v}$ are unit eigenvectors corresponding to different eigenvalues, and $|\epsilon|>0$ is small. Furthermore, from (14), we obtain $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}=0$ if and only if $\mathbf{y}$ is a unit eigenvector for $M$. However, this implies that $\mathbf{y}$ is fixed by Proposition 3. That is, $\mathbf{y}$ cannot exhibit strong cospectrality, a contradiction to the fact that $\mathbf{y}$ is involved in PST. Therefore, $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}<0$.
@@
Next we seek the minimum value of $(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}$ subject to the constraint that $\sum_{j=1}^{\ell}a_{j}=1$ and $a_{j}\geq 0$ for all $j=1,\ldots,\ell.$ Suppose that we have distinct indices $j_{1},j_{2}$ such that $a_{j_{1}},a_{j_{2}}>0.$ For each $j=1,\ldots,\ell$, consider the coefficients $b_{j}$ given by $b_{j}=a_{j}$ whenever $j\neq j_{1},j_{2},b_{j_{1}}=a_{j_{1}}+\epsilon$, and $b_{j_{2}}=a_{j_{2}}-\epsilon,$ where $\epsilon$ is sufficiently small. It is straightforward to show that
$\begin{split}(\sum_{j=1}^{\ell}b_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}&=(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}+\epsilon(\lambda_{j_{1}}-\lambda_{j_{2}})(2\sum_{j=1}^{\ell}a_{j}\lambda_{j}-(\lambda_{j_{1}}+\lambda_{j_{2}}))+\epsilon^{2}(\lambda_{j_{1}}-\lambda_{j_{2}})^{2}.\end{split}$ (15)
@@
Observe that if there is a third index $j_{3}\neq j_{1},j_{2}$ such that $a_{j_{3}}>0,$ then because $\lambda_{j_{2}}\neq\lambda_{j_{3}}$, we get that one of $2\sum_{j=1}^{\ell}a_{j}\lambda_{j}-(\lambda_{j_{1}}+\lambda_{j_{2}})$ and $2\sum_{j=1}^{\ell}a_{j}\lambda_{j}-(\lambda_{j_{1}}+\lambda_{j_{3}})$ is nonzero. In particular, if $2\sum_{j=1}^{\ell}a_{j}\lambda_{j}-(\lambda_{j_{1}}+\lambda_{j_{2}})\neq 0$, then by virtue of (15), we can choose a value of $\epsilon$ so that
$(\sum_{j=1}^{\ell}b_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}b_{j}\lambda_{j}^{2}<(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}.$
In this case, $(\sum_{j=1}^{\ell}a_{j}\lambda_{j})^{2}-\sum_{j=1}^{\ell}a_{j}\lambda_{j}^{2}$ cannot attain the minimum value. So in order to find the minimum value, it suffices to focus on expressions of the form $(a\lambda_{j_{1}}+(1-a)\lambda_{j_{2}})^{2}-(a\lambda_{j_{1}}^{2}+(1-a)\lambda_{j_{2}}^{2})$ where $a\in[0,1]$ and $\lambda_{j_{1}},\lambda_{j_{2}}\in\{\lambda_{1},\ldots,\lambda_{\ell}\}.$ Let $\lambda_{\max},\lambda_{\min}$ denote the largest and smallest elements of $\{\lambda_{1},\ldots,\lambda_{\ell}\},$ respectively. Since
$(a\lambda_{j_{1}}+(1-a)\lambda_{j_{2}})^{2}-(a\lambda_{j_{1}}^{2}+(1-a)\lambda_{j_{2}}^{2})=-a(1-a)(\lambda_{j_{1}}-\lambda_{j_{2}})^{2},$
we find that the minimum value for $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}$ is given by $-\frac{1}{2}(\lambda_{\max}-\lambda_{\min})^{2}$, which is attained by $\mathbf{x}=\frac{1}{\sqrt{2}}(\mathbf{u}+\mathbf{v})$ and $\mathbf{y}=\frac{1}{\sqrt{2}}(\mathbf{u}-\mathbf{v}),$ where $\mathbf{u},\mathbf{v}$ are unit eigenvectors corresponding to $\lambda_{\max}$ and $\lambda_{\min},$ respectively. We summarize the above discussion as follows.
@@
Suppose perfect state transfer occurs between unit vectors $\mathbf{x}$ and $\mathbf{y}$ at time $\tau$ relative to $M$. Let $\lambda_{\max}=\max\sigma_{\mathbf{x}}(M),\lambda_{\min}=\min\sigma_{\mathbf{x}}(M).$ Then
$0>\ \frac{d^{2}f}{dt^{2}}\bigg{|}_{\tau}\ \geq\ -\frac{1}{2}(\lambda_{\max}-\lambda_{\min})^{2}.$
@@
Consider the Petersen graph with adjacency matrix $A,$ which has eigenvalues $3,1,-2$. Let $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$ be unit eigenvectors corresponding to $3,1$ and $-2,$ respectively and form $\mathbf{x}\in\mathbb{R}^{10}$ as $\mathbf{x}=c_{1}\mathbf{v}_{1}+c_{2}\mathbf{v}_{2}+c_{3}\mathbf{v}_{3}$ where $c_{1},c_{2},c_{3}\in\mathbb{R}\backslash\{0\},c_{1}^{2}+c_{2}^{2}+c_{3}^{2}=1.$ Setting $\mathbf{y}=-c_{1}\mathbf{v}_{1}-c_{2}\mathbf{v}_{2}+c_{3}\mathbf{v}_{3}$, we find that there is PST from $\mathbf{x}\mathbf{x}^{\top}$ to $\mathbf{y}\mathbf{y}^{\top}$ at time $\pi.$ According to Theorem 61, we have $0>\frac{d^{2}f}{dt^{2}}\big{|}_{\pi}\geq-\frac{25}{2}$.
Combining Theorem 61 with the proof of Corollary 59, we obtain the connected unweighted $n$-vertex graphs and the unit vectors admitting PST that attain the least value of $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}$. These unit vectors are precisely those that minimize the PST time amongst all connected unweighted $n$-vertex graphs.
Corollary 63.
Amongst all connected unweighted $n$-vertex graphs that admit perfect state transfer between unit vectors at time $\tau$ relative to $L$ (respectively, relative to $A$), the least value of $\frac{d^{2}f}{dt^{2}}\big{|}_{\tau}$ is attained by the join graph in Theorem 58(1) (respectively, Theorem 58(2)) and unit vectors $\mathbf{x}$ and $\mathbf{y}$, where $\mathbf{y}$ is given in Corollary 59(1) (respectively, Corollary 59(2)).
Acknowledgement. C. Godsil is supported by NSERC grant no. RGPIN-9439. S. Kirkland is supported by NSERC grant no. RGPIN-2025-05547. H. Monterde is supported by the University of Manitoba Faculty of Science and Faculty of Graduate Studies. We thank the Department of Mathematics and the Graduate Mathematics Society at the University of Manitoba for supporting the research visit of C. Godsil, where we all started to work on this project.